# Internal-Coordinate Density Modelling of Protein Structure: Covariance Matters

### 07 Dec 2023 v2

@@ 1 Introduction
In this paper, we investigate whether density modelling of full-size protein backbones in internal coordinates is feasible. We empirically demonstrate the difficulty in estimating the covariance structure of internal coordinates from data, and instead propose a technique for inducing the covariance structure by imposing constraints on downstream atom movement using the Lagrange formalism. Rather than estimating the covariance structure from scratch, we can instead modulate the covariance structure by choosing appropriate values for allowed fluctuations of downstream atoms. We demonstrate the procedure in the context of a variational autoencoder (Fig. 1).
@@ 3.2 Lagrange formalism to incorporate constraints
$\tilde{\mathcal{D}}(p,q)=$ $\int p(\Delta\bm{\kappa})\ln\frac{p(\Delta\bm{\kappa})}{q(\Delta\bm{\kappa})}\mathrm{d}\Delta\bm{\kappa}+\lambda_{0}\left(\int p(\Delta\bm{\kappa})\mathrm{d}\Delta\bm{\kappa}-1\right)$
$+\sum_{m=1}^{M}\lambda_{m}\left(\int\Delta x_{m}^{2}p(\Delta\bm{\kappa})\mathrm{d}\Delta\bm{\kappa}-C_{m}\right).$ (5)
Next, we take the functional derivative and set it to zero: $\frac{\partial\tilde{\mathcal{D}}(p,q)}{\partial p(\Delta\bm{\kappa})}=0$, leading to the following well-established result (Jaynes, 1957; Kesavan & Kapur, 1989) [^2] Note that $\frac{\partial}{\partial y}\ln{y/q}+\lambda_{0}(y-1)+\sum_{m=1}^{M}\lambda_{m}\left(\Delta x_{m}^{2}y-C_{m}\right)=\ln{\frac{y}{q}}+y\cdot\frac{1}{y}+\lambda_{0}+\sum_{m=1}^{M}\lambda_{m}\Delta x^{2}_{m}$:
@@ 4.1 Test cases
We also include two test cases for larger fluctuations following a multimodal distribution in a high data regime. Both are known as “fast-folders” and the MD datasets were obtained from Lindorff-Larsen et al. (2011). We refer the reader to this work for detailed descriptions of the simulations. Chignolin (cln025) is a peptide with a hairpin structure, containing 10 residues and thus 30 backbone atoms. The simulation is 106 $\upmu$s long, saved at a 200 ps interval, resulting in 534.743 data points. The second test case, 2f4k, is the chicken villin headpiece, with 35 residues and 105 backbone atoms. The simulated trajectory is 125 $\upmu$s and also saved every 200 ps, yielding 629.907 structures.
@@ 4.2 Metrics
For the unimodal setting, we choose two simple measures of local and global structure, respectively. To evaluate local structure fluctuations, we show Ramachandran plots, a well-known visualization tool in the context of protein structures, where $\phi$ and $\psi$ dihedrals, which are the torsional angles around the $N-C_{\alpha}$ and $C_{\alpha}-C$ bonds, are plotted against each other. As a global measure, we report the variance over atom positions, averaged over three dimensions, across superposed (i.e. structurally aligned) samples to evaluate global structure fluctuations.
@@
Apart from comparing the generated samples from our model to reference distributions from MD or NMR, we include four baselines. The first baseline, named “$\kappa$-prior (fixed)” is a VAE trained to predict $/mu_{\kappa}$ given a fixed covariance matrix that is equal to $\bm{\Sigma}^{-1}_{\rm{\bm{\kappa},prior}}$. In other words, this is the same as our full VAE setup, but omitting the imposed 3D constraints. The second baseline is “$\kappa$-prior (learned)”, which corresponds to a more standard VAE-setting where the decoder directly outputs a mean and a variance (i.e. a diagonal covariance matrix). The third baseline does not involve a VAE, but samples structures from a multivariate Gaussian with a mean based on the dataset and a precision matrix computed by a standard estimator. This is an emperical estimator for MD datasets, and an Oracle Approximating Shrinkage (OAS) estimator (Chen et al., 2010) for NMR datasets, since emperical estimators do not converge for such low amounts of samples. Finally, we include the flow-based model from Köhler et al. (2023) as a fourth baseline, which, to the best of our knowledge, is the current state of the art in density modelling for internal coordinate representations.
@@ 4.3 Internal-coordinate density modelling results
The unimodal, low data regime test cases exhibit small fluctuations around the native protein structure, where the largest fluctuations correspond to the loops connecting different secondary structure elements. Fig. 4 demonstrates that 1unc, 1fsd and 1pga structures sampled from the VAE conform to local and global constraints, with valid Ramachandran distributions when compared to the reference as well as improved atom position variance along the chain compared to the baselines (see quantitative results in Table A2). Even in the extremely low data regime of 25 and 41 data points for 1unc and 1fsd, respectively (top two rows in Fig. 4), the VAE is able to estimate a full covariance matrix that approximates the distribution better than the baselines, especially in loop regions. This can lead to unphysical structures with backbone crossings, even though the local structure is preserved. These effects can also be observed in the 3D visualization of the sampled ensembles in Fig. A2.
@@
The third test case, 1pga (bottom row in Fig. 4), has a more complex structure with two $\upbeta$-strands at the $\mathrm{N}$-terminus forming a sheet together with two $\upbeta$-strands from the $\mathrm{C}$-terminus. These global constraints are not captured well by the baselines in this low data regime, resulting in very high fluctuations in loop regions which violate the native structure (additional visualizations can be found in Fig. A2). For our VAE, we see the benefits of imposing global constraints, resulting in much better density estimation compared to the baselines. Moreover, we can control the interplay between local and global constraints by adjusting the hyperparameters of our model, as exemplified in Section D.1. However, the complexity of this protein prevents perfect density estimation in a low data regime. Interestingly, we show in Section C.1 that the variance of the atom positions highly correlates to the imposed constraints $C$ calculated from a set of predicted Lagrange multipliers using Section 3.3.
@@ Appendix A Experimental details
We use a standard U-Net (Ronneberger et al., 2015) to predict atom fluctuation constraints $\lambda$ from the mean predictions that the decoder outputs. The predicted mean for the internal degrees of freedom $\bm{\mu_{\kappa}}$ is translated into a mean structure $\bm{\mu_{x}}$ using pNeRF (AlQuraishi, 2018), from which we calculate a pairwise distance matrix. This $M\times M$ matrix with a single “channel” serves as the input to the U-Net, which scales the number of channels up to $1024$ in four steps before scaling back down to one channel in four steps.
@@ D.2 Multimodal
When sampling from the VAE, where constraints are imposed on top of the $\kappa$-prior, there is interplay between $a$ and $w_{\mathrm{aux}}$, as shown in Fig. A6. Even though the exact trend is less clear here, the relative values of the hyperparameters have an observable influence on e.g. the width of the ”bridge” between the topmost two clusters, the size of the higher-energy downward extrusion of the top left cluster, and the spread towards the less populated cluster on the bottom right.

### 24 Jan 2024 v3

@@ 1 Introduction
In this paper, we investigate whether density modelling of full-size protein backbones in internal coordinates is feasible. We empirically demonstrate the difficulty in estimating the covariance structure of internal coordinates from data, and instead propose a technique for inducing the covariance structure by imposing constraints on downstream atom movement using the Lagrange formalism. Rather than estimating the covariance structure from scratch, we can instead modulate the covariance structure by choosing appropriate values for allowed fluctuations of downstream atoms (Fig. 1). We demonstrate the procedure in the context of a variational autoencoder.
@@ 3.2 Lagrange formalism to incorporate constraints
$\tilde{\mathcal{D}}(p|q)=\int p(\Delta\bm{\kappa})\ln\frac{p(\Delta\bm{\kappa})}{q(\Delta\bm{\kappa})}\mathrm{d}\Delta\bm{\kappa}+\lambda_{0}\left(\int p(\Delta\bm{\kappa})\mathrm{d}\Delta\bm{\kappa}-1\right)+\sum_{m=1}^{M}\lambda_{m}\left(\int\Delta x_{m}^{2}p(\Delta\bm{\kappa})\mathrm{d}\Delta\bm{\kappa}-C_{m}\right).$
Next, we take the functional derivative and set it to zero: $\frac{\partial\tilde{\mathcal{D}}(p,q)}{\partial p(\Delta\bm{\kappa})}=0$, leading to the following well-established result (Jaynes, 1957; Kesavan & Kapur, 1989) [^2] Note that $\frac{\partial}{\partial y}\left(y\ln{y/q}+\lambda_{0}(y-1)+\sum_{m=1}^{M}\lambda_{m}\left(\Delta x_{m}^{2}y-C_{m}\right)\right)=\ln{\frac{y}{q}}+y\cdot\frac{1}{y}+\lambda_{0}+\sum_{m=1}^{M}\lambda_{m}\Delta x^{2}_{m}$:
@@ 4.1 Test cases
We also include two test cases for larger fluctuations following a multimodal distribution in a high data regime. Both are known as “fast-folders” and the MD datasets are available upon request from the authors of Lindorff-Larsen et al. (2011). We refer the reader to this work for detailed descriptions of the simulations. Chignolin (cln025) is a peptide with a hairpin structure, containing 10 residues and thus 30 backbone atoms. The simulation is 106 $\upmu$s long, saved at a 200 ps interval, resulting in 534.743 data points. The second test case, 2f4k, is the chicken villin headpiece, with 35 residues and 105 backbone atoms. The simulated trajectory is 125 $\upmu$s and also saved every 200 ps, yielding 629.907 structures.
@@ 4.2 Metrics
For the unimodal setting, we choose two simple measures of local and global structure, respectively. To evaluate local structure fluctuations, we show Ramachandran plots, a well-known visualization tool in the context of protein structures, where $\phi$ and $\psi$ dihedrals, which are the torsional angles around the $N-C_{\alpha}$ and $C_{\alpha}-C$ bonds, are plotted against each other. As a global measure to evaluate the overall 3D structural fluctuations, we report the variance over atom positions, averaged over three dimensions, across superposed (i.e. structurally aligned) samples.
@@
Apart from comparing the generated samples from our model to reference distributions from MD or NMR, we include four baselines. The first baseline, named “$\kappa$-prior (fixed)” is a VAE trained to predict $\bm{\mu_{\kappa}}$ given a fixed covariance matrix that is equal to $\bm{\Sigma}^{-1}_{\rm{\bm{\kappa},prior}}$. In other words, this is the same as our full VAE setup, but omitting the imposed 3D constraints. The second baseline is “$\kappa$-prior (learned)”, which corresponds to a more standard VAE-setting where the decoder directly outputs a mean and a variance (i.e. a diagonal covariance matrix). The third baseline does not involve a VAE, but samples structures from a multivariate Gaussian with a mean based on the dataset and a precision matrix computed by a standard estimator. This is an emperical estimator for MD datasets, and an Oracle Approximating Shrinkage (OAS) estimator (Chen et al., 2010) for NMR datasets, since emperical estimators do not converge for such low amounts of samples. Finally, we include the flow-based model from Köhler et al. (2023) as a fourth baseline, which, to the best of our knowledge, is the current state of the art in density modelling for internal coordinate representations.
@@ 4.3 Internal-coordinate density modelling results
The unimodal, low data regime test cases exhibit small fluctuations around the native protein structure, where the largest fluctuations correspond to the loops connecting different secondary structure elements. Fig. 4 demonstrates that 1unc, 1fsd and 1pga structures sampled from the VAE conform to local and global constraints, with valid Ramachandran distributions when compared to the reference as well as improved atom position variance along the chain compared to the baselines (see quantitative results in Table A2). Even in the extremely low data regime of 25 and 41 data points for 1unc and 1fsd, respectively (top two rows in Fig. 4), the VAE is able to estimate a full covariance matrix that approximates the distribution better than the baselines, especially in loop regions. Consequently, the baselines show a higher tendency to sample unphysical structures with backbone crossings, even though the local structure is preserved. These effects can also be observed in the 3D visualization of the sampled ensembles in Fig. A2.
@@
The third test case, 1pga (bottom row in Fig. 4), has a more complex structure with two $\upbeta$-strands at the $\mathrm{N}$-terminus forming a sheet together with two $\upbeta$-strands from the $\mathrm{C}$-terminus. These global constraints are not captured well by the baselines in this low data regime (400 data points), resulting in very high fluctuations in loop regions which violate the native structure (additional visualizations can be found in Fig. A2). For our VAE, we see the benefits of imposing global constraints, resulting in much better density estimation compared to the baselines. Moreover, we can control the interplay between local and global constraints by adjusting the hyperparameters of our model, as exemplified in Section D.1. However, the complexity of this protein prevents perfect density estimation in a low data regime. Interestingly, we show in Section C.1 that the variance of the atom positions highly correlates to the imposed constraints $C$ calculated from a set of predicted Lagrange multipliers using Section 3.3.
@@ 6 Discussion
7 Code and data availability
Code and in-house generated data will be made available upon acceptance. A minimal version of the code base is added as supplementary material.
8 Acknowledgements
The work was supported by the Novo Nordisk Foundation (project grant nr NNF18OC0052719) and conducted within the Center for Basic Machine Learning Research in Life Science (MLLS, grant nr NNF20OC0062606).
@@ Appendix A Experimental details
We use a standard U-Net (Ronneberger et al., 2015) to predict atom fluctuation constraints $\lambda$ from the mean predictions that the decoder outputs. The predicted mean for the internal degrees of freedom $\bm{\mu_{\kappa}}$ is translated into a mean structure $\bm{\mu_{x}}$ using pNeRF (AlQuraishi, 2018), from which we calculate a pairwise distance matrix. This $M\times M$ matrix with a single channel serves as the input to the U-Net, which scales the number of channels up to $1024$ in four steps before scaling back down to one channel in four steps.
@@ D.2 Multimodal
When sampling from the VAE, where constraints are imposed on top of the $\kappa$-prior, there is interplay between $a$ and $w_{\mathrm{aux}}$, as shown in Fig. A6. Even though the exact trend is less clear here, the relative values of the hyperparameters have an observable influence on e.g. the width of the "bridge" between the topmost two clusters, the size of the higher-energy downward extrusion of the top left cluster, and the spread towards the less populated cluster on the bottom right.